x+y=k
x-y=k
the solution to the system shown is
(0,k)
(2,2)
(k,0)

Answer by stanbon(75887)

x+y=k
x-y=k
-----
Add and solve for "x":
2x = 2k
x = k
----
Solve for "y"
k + y =k
y = 0
-----
Solution: (k,0)
Cheers,
Stan H.
---------------

the solution to the system shown is
(0,k)
(2,2)
(k,0)